Status Report

NASA Space Station On-Orbit Status 2 August 2004

By SpaceRef Editor
August 2, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 2 August 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.  Underway:  Week 14 for Expedition 9.

Crew activities today focused on final preps for tonight’s Orlan spacewalk (EVA-10).  Wakeup was slipped by 30 min. to 2:30am EDT.  Sleep cycle began only 6.5 hrs later, at 9:00am.  Second wakeup is tonight at 5:30pm, followed by a delayed sleep time at 3:00pm tomorrow, 8/3.

As part of morning station inspection, CDR Padalka today did the regular (monthly) routine checkup in the DC-1 “Pirs” docking module on the AZS circuit breakers on the BVP Amp Switch Panel — they should all be On — and the LEDs (light-emitting diodes) of the 14 fuses in Fuse Panels BPP-30 & BPP-36 (last time done: 7/1).

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

FE/SO Fincke meanwhile performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).

Fincke then transferred SODF (station operations data file) books and the CCPK (Crew Contamination Protection Kit) from the Lab to the DC1 docking compartment.   [The hardcopy transfer involved the POC and Medical Checklist books, CD library copies, plus verification that the two emergency (EMER) books are available in the Russian segment (RS).  The CCPK’s purpose is to provide protection inside the station against the unlikely event of crew contamination from toxic thruster plume residue picked up by during the spacewalk.]

Both crewmembers made final preparations on their Orlan-M suits and EVA tools/equipment to be taken out tonight.  These preps included

  • Filling & installing the DIDBs (disposable in-suit drink bags), postponed from 7/29;
  • Equipping both Orlans (in pocket on left calf) with a “Pille-MKS” radiation sensor.   [These sensors, A0307 &A0309, were removed from their exposure locations in the RS after recording their dose measurements; a third sensor, A0310, was placed in the SM cabin for background readings.   [“Pille” has ten sensors normally situated at various locations in the RS (port cabin window, stbd cabin window, ASU toilet facility, control panel, etc.).  Dosage values are called down or downlinked via Regul Paket/Email or OCA.], and
  • Installing freshly charged NiMH (nickel metal hydride) batteries in the EHIP (EMU helmet interchangeable portable) lights, borrowed from U.S. EMUs and jerry-rigged on the Orlans.

Other station preparations by Mike Fincke for the 6-hr. spacewalk included:

  • Configuring the EPS (electrical power system) as required for USOS/RS  (U.S. segment/Russian segment) module isolation for EVA;
  • Reconnecting the UOP (utility outlet panel) bypass power cable at both Lab and Cupola RWS DCPs (robotics workstation/display & control panels);
  • Setting up two Sony PD100 camcorders in the Lab and Node for situational awareness during the spacewalk; and
  • Configuring the Ethernet OpsLAN (Operations Local Area Network) for the RS/USOS hatch closure, and subsequent reopening after the EVA.

Gennady meanwhile shuttered SM windows #6 & #8 with their external covers to protect them from accidental contact during the spacewalk on the RS.

The CDR also disconnected the new Sputnik-SM “Kenwood D700” amateur radio station to protect the spacewalkers from inadvertent antenna RF emission.

Padalka completed the daily routine inspection of the SM’s SOZh life support system (including ASU toilet facilities), while Fincke conducted the regular routine status checkup of the autonomous ISS-7 Lab payloads (PCG-STES010, SAMS, MAMS).

Shortly before sleep time (9:00am), the CDR once more broke out the “Urolux” equipment, setting it up for the Russian biochemical urine test (PZE MO-9), a standard requirement before and after Orlan-suited activities (in this case post-EVA).   [The MO-9 analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.]

>>> EVA-10 Look-ahead:   Hatch closure between SM PkhO (Transfer Compartment) and DC1 is set for early tomorrow morning (8/3) at  ~1:30am EDT, followed by Orlan ingress (1:40am) and start of 30-min prebreathe at ~2:22am.  DC1 will then be depressurized, first to 15 mmHg (Torr), for subsequent leak checking, then to full vacuum and EV 1 hatch opening at ~3:14am.  Unlike EVA-9, the ~6-hr. spacewalk tonight takes place on the RS (DC1 and SM).  Its objectives include (a) installation of protective devices on DC1 hatch circular handrail brackets and gap spanners between existing DC1 handrails, (b) still photography and replacement of a collection tray of the Kromka-3 contamination experiment and of a removable materials sample cassette (SKK) on the SM, (c) replacement of laser retro reflectors (LRRs) for ATV (automated transfer vehicle) docking with improved LRRs, (d) installation of an internal Visual Video Target (VVT) and two antennas (WAS1 & 2) for the Proximity Communication Equipment (PCE) with their cabling, (e) removal of a Platan-M detector unit, etc.  The spacewalk is nominally expected to last ~5h 42m, with hatch closing at ~8:56am.  DC1 will then be repressed from the SM cabin air, first to 260 mmHg for leak checks, then to final 752 mmHg.  Orlan-M suits will be doffed at ~9:27am and hatch to SM opened at ~9:45am tomorrow morning.

The station continues to fly in LVLH XVV attitude (local vertical/local horizontal — X-axis in velocity vector, i.e. bow forward), until tomorrow morning (00:45am), shortly before the EVA-10.

Major upcoming events:

  • Orlan EVA-10 from DC-1 — 8/3 (hatch open: 3:14am EDT);
  • EVA debrief, etc. — 8/4;
  • Progress 15P launch — 8/11 (1:01am EDT);
  • Progress 15P docking — 8/14 (2:05am EDT).

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Manila, Philippines (nadir pass), and Lightning, Philippines (area of general storminess may hold up for lightning opportunities).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 4:32am EDT [= epoch]):

  • Mean altitude — 357.5 km
  • Apogee height — 361.7 km
  • Perigee height — 353.4km
  • Period — 91.7 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0006148
  • Solar Beta Angle — -18.3 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32556

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.